OpenStudy (anonymous):

How do you do this: 7 (y-2) = -y + 8y - 14 2(4b - 3) = 8 - 6

OpenStudy (mertsj):

Remove the parenthesis first by using the distributive property

OpenStudy (anonymous):

is this rite 7(y-2)= -y+8y-14 7y-14=7y-14 y= 0 2(4b-3)= 8-6 8b-24= 2 8b= 26 b= 13/4

OpenStudy (mertsj):

Actually the first one would be 0=0 which means that any number at all would make the sentence true.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thank you :D

OpenStudy (mertsj):